The motorboat traveled 3 times as fast as the sloop. In fact, the motorboat traveled 48 miles in 2 hours less than it took the sloop to travel 24 miles. How fast did each of them travel? What were the times of each?
Answer by lwsshak3(11628)

The motorboat traveled 3 times as fast as the sloop. In fact, the motorboat traveled 48 miles in 2 hours less than it took the sloop to travel 24 miles. How fast did each of them travel? What were the times of each?
**
let x=speed of sloop
3x=speed of motorboat
travel time=distance/speed
..
48/3x=24/x-2
LCD: 3x
48=72-6x
6x=24
x=4
3x=12
..
48/3x=4
24/x=6
..
speed of sloop=4 mph
speed of motorboat=12 mph
travel time of sloop=6 hrs
travel time of motorboat=4 hrs